OpenStudy (anonymous):

FAN AND MEDAL Solve this equation step by step by using like terms -6 1/3 - 1/2 ( 1/2 + y )

OpenStudy (anonymous):

@campbell_st @aaronq @Luigi0210 @iGreen @.Sam. @SithsAndGiggles @confluxepic @TheSmartOne @Nnesha @undeadknight26 @pooja195 @bibby

OpenStudy (kendricklamar2014):

Can you draw the problem @stylestar5678

OpenStudy (anonymous):

yeah just give me a second

OpenStudy (kendricklamar2014):

\[Ok:)\]

OpenStudy (anonymous):

Actually I cant. It doesn't let me on the iPad? Should I retype it with the equation thingy?

OpenStudy (kendricklamar2014):

Yes Please

OpenStudy (anonymous):

\[6\frac{ 1 }{ 3 }-\frac{ 1 }{ 2}(\frac{ 1 }{ 2 }+y)\]

OpenStudy (kendricklamar2014):

* -6

OpenStudy (anonymous):

oh yeah sorry i forgot to add the negative

OpenStudy (kendricklamar2014):

First we distribute:\[-6(1/3) + -1/2y + -1/4\]

OpenStudy (kendricklamar2014):

Which equals: 

OpenStudy (kendricklamar2014):

Now you Combine Like Terms which equals: